OpenStudy (anonymous):

Find p,q, and r if f(0)=-5, f(3)=-5, and f(5)=-35 for f(x)=px^2+qx+r, show the process please

OpenStudy (anonymous):

for f(0) -5 = p (0^2) + q(0) + r -5 = r Do you understand this?

OpenStudy (anonymous):

yes, go on

OpenStudy (anonymous):

use the same concept for the others.

OpenStudy (anonymous):

that doesn't give you a complete answer for the others

OpenStudy (anonymous):

I'm not gonna help you if you are just looking for answers. Put in some effort.

OpenStudy (anonymous):

By using the same concept, I will get -5=9p+3q+r, and -35=25p+5q+r

OpenStudy (anonymous):

Which isn't complete

OpenStudy (anonymous):

There are more steps afterward, I am just unsure of them

OpenStudy (anonymous):

as earlier, remember i found a value for r?

OpenStudy (anonymous):

Yeah, by plugging in zero, but that doesnt work for the other functions

OpenStudy (anonymous):

nos substitute that r value in the other two equations.

OpenStudy (anonymous):

okay so then I'd have -5=9p+3q-5 and -35=25p+5q-5. And then I use either substitution or elimination to solve the equations right?

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

okay one sec lemme work it out and see if i get the right answer

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Okay so the answer is -5=r, -3=p, and 9=q right??? @mickey4691

OpenStudy (anonymous):

verify the signs for q and p.

OpenStudy (anonymous):

nuuuuuu I got it wrong???? :(

OpenStudy (anonymous):

darn.

OpenStudy (anonymous):

the signs.

OpenStudy (anonymous):

-5=9p+3q-5 that means 9p+3q= 0 then q= -3p

OpenStudy (anonymous):

-35=25p+5q-5 then 25p + 5 q = 30 then 5p + q = 6 then 5p - 3p = 6 2p=6 p = 3

OpenStudy (anonymous):

once p = 3

OpenStudy (anonymous):

q = -3p

OpenStudy (anonymous):

q= -3 (3) = -9

OpenStudy (anonymous):

wait but 25p+5q= -30 not +30

OpenStudy (anonymous):

good job. then you are correct lol.

OpenStudy (anonymous):

You deserve the medal lol

OpenStudy (anonymous):

i was only testing you lol. i'm kidding.

OpenStudy (anonymous):

oh okay thank you so much I wouldn't have gotten there without you :)